OpenStudy (anonymous):

3. find y if the line through A(-1,2) and B(3,5) is parallel to the line through (0,-1) and (6,y)

OpenStudy (anonymous):

Help!!!

OpenStudy (anonymous):

find gradient of AB then equate it to gradient of the other line (0.-1) and (6,y)

OpenStudy (phi):

what is the slope (change in Y divided by change in X) between A(-1,2) and B(3,5) ?

OpenStudy (anonymous):

Parellel, means the "gradient" of both the slopes are equal. FInd the gradient of line AB \[\frac{y_{2}-y_{1}}{x_{2}-x_{1}} = m\] equate m into \[\frac{y-(-1)}{6-0} = m\] SOlve it u get Y

OpenStudy (anonymous):

is y=3/4???

OpenStudy (phi):

what did you get for the slope between A and B ?

OpenStudy (anonymous):

3/4 is the gradient

OpenStudy (anonymous):

to find it substitute gradient for the the second slope

OpenStudy (anonymous):

sory, my mistake, the slope is 3/4

OpenStudy (anonymous):

should i cross multiply??

OpenStudy (anonymous):

yes u have to

OpenStudy (phi):

you could cross multiply, but I would just multiply both sides by 6

OpenStudy (anonymous):

why 6???

OpenStudy (anonymous):

\[y-(-1) = 6m\]

OpenStudy (anonymous):

where is that equation from??

OpenStudy (phi):

the 2nd equation, posted by thus, is the slope of the 2nd line

OpenStudy (anonymous):

Do you know the equation to find the gradient of the slope, when the coordinates of 2 points are given?

OpenStudy (anonymous):

shi*,, i think i'm just gonna cross multiply them. xD

OpenStudy (anonymous):

so what did u get for Y?

OpenStudy (anonymous):

UHM.. -22/4

OpenStudy (anonymous):

i think u got your mathematics wrong. \[y-(-1) = 6m..... m = 3/4\] \[y + 1 = 6\times \frac{3}{4}\] \[y = \frac{9}{2} - 1\] \[y = \frac{7}{2}\]